OpenStudy (anonymous):

Let f(x) = ax^2 and g(x) = bx + c, where a, b, and c are constants. Compute f (g(x)) and g (f(x)). Determine for which constants a, b, and c it is true that f ( g(x)) = g(f(x))

OpenStudy (anonymous):

I know how to get f(g(x)) I got a(bx+c)^2 and g(f(x))=b(ax^2)+c but when I made them equal

OpenStudy (anonymous):

I can't solve for anything

OpenStudy (anonymous):

expanding to get b^2x^2+2bx +c^2=abx^2+c if they are equal

OpenStudy (anonymous):

so b^2=ab b^2-ab=0 b(b-a)= 0 b=0 b=a

OpenStudy (anonymous):

oops made an error

OpenStudy (anonymous):

a(b^2x^2+2bx +c^2)=abx^2+c

OpenStudy (anonymous):

ab^2=ab ab^2-ab=0 ab(b-1) = 0 so b = 1

OpenStudy (anonymous):

sorry I don't see how you got that ab^2=ab from the equation?

OpenStudy (anonymous):

with the x terms 2ab =0 hmm with the constants ac^2 = c ac^2 - c = 0 c(ac -1)= 0

OpenStudy (anonymous):

I equated the coefficients

OpenStudy (anonymous):

equate the coefficients on the left from f(g(x)) and g(f(x)) ... I used what u said u got

OpenStudy (anonymous):

oh so then 2abx=0 and c^a=c?

OpenStudy (anonymous):

oops c^2a=c

OpenStudy (anonymous):

if they are equal then the coefficients of x^2, x and the constants must be equal ... hope that helps

OpenStudy (anonymous):

That make sense then you just leave them like that? you can't really get actually values